Let n be an even number. How many ternary strings (i.e. strings over the alphabet 10, 1,2]) of length n are there in which the only places that zeroes can appear are in the odd-numbered positions?
Answer #1

Only 1 and 2 can be placed at even positions, so number of combinations for even positions = n/2*2...........a

and 0,1,2, can be placed at odd positions, so no. of combinations for odd positions= n/2*3............................b

Therefore total no. of combinations= a*b

=n*n*2/3